# Interface-sensitive microwave loss in superconducting tantalum films sputtered on c-plane sapphire

### v1

@@ III.3 Modifying the Ta/ interface
For the second test, referred to as Sample C, the Al2O3(0001) substrate was treated with an argon (Ar) plasma in the growth chamber in-situ at room temperature before the high vacuum outgas and growth of Ta directly on the substrate at 630. The plasma exposure was performed at 10mTorr using an RF power of 50W for 120 seconds. This plasma condition was found to etch aluminum at 1.6 nm/min in a separate calibration that is provided here for repeatability. The outgas temperature (880) used is not expected to result in restructuring of the Al2O3(0001) surface, which occurs at temperatures above 1000Wildes, Mayer, and Theis-Brol (2001).
Ar plasma exposure is expected to both etch the sapphire substrate and induce structural damage to the sapphire surface. The structural damage to the sample surface was observed using in-situ RHEED, where the bright diffraction pattern of the sapphire observed initially became dim (though still observable) after plasma exposure, and was accompanied by an increase in the diffuse background, indicating a less ordered surface. RHEED patterns of this sample after Ta growth were consistent with a textured film having both out-of-plane and in-plane texture. This texture was confirmed with XRD after the sample was cooled and removed from the growth chamber.
@@
Figure 4(b) shows higher resolution scans of the -Ta(222) reflections from which the difference in the out-of-plane lattice constant of the Ta films grown with and without the Nb inter-layer is computed to be less than 1%. Fig. 4(c) shows - scans (in-plane sample rotation) of non-specular substrate and film peaks of Sample A, which confirm that the Ta is epitaxially arranged to the substrate having the same rotational relationship stated earlier for the epitaxial Nb. Non-specular XRD measurements of Sample C (not shown) were also performed which showed in-plane texturing with 6 rotational domains.
Figure 5 includes atomic force micrographs taken from Samples A, B, and C. For Samples A and B, (the epitaxial Ta with and without the 5nm thick Nb interlayer), the surface morphology was smooth and shapeless with little difference between the two. For Sample C (the Ta film grown on plasma-treated substrate), elongated, grain-like features were observed, consistent with previous reports of textured Ta and Nb films Clavero et al. (2012); Jones et al. (2023). The direction of the elongation reflects the in-plane texturing of the film where the same 6 rotational domains measured with XRD are observed. RMS roughness values extracted from AFM data is included in Table 2 for all samples.
Figure 6 shows temperature dependent resistivity measurements of Samples A, B, and C. For the epi--Ta(111) (Samples A and B), the 5nm Nb interlayer had little effect on film resistance, RRR, or . The RRR of the textured Ta(110) film grown on the plasma damaged sapphire (Sample C) was lower, reflecting greater disorder in the polycrystalline film. The of all -Ta films was between 4.2-4.5K and the complete data is included in Table 2.
@@ IV Discussion
Consider Samples A, B, and C as indicated in Table 2. All three films were grown on c-plane sapphire at the same temperature of 630: Sample A is epi--Ta(111) grown directly on sapphire, Sample B is epi--Ta(111) grown on a 5nm thick Nb interlayer on sapphire, and Sample C is textured--Ta(110) grown on plamsa treated sapphire. The of resonators made from Sample A are markedly lower than B and C. The microscopic mechanism of this loss is of interest, and some potential sources are discussed here in light of consistency with the results presented in this work.
@@
Initially, the origin of the observed interface loss was thought to be either due to reactions between the Ta and the Al2O3 substrate or chemical contamination on the sapphire surface before growth, but neither explanation is fully consistent with the results. Fist addressing interface contamination: the substrates used for all samples presented in this study (with exception of Sample C) were prepared in the same way and thus had very similar levels of surface contamination before film growth. High was observed for all Nb films grown, and also for the Ta films grown at lower temperature. Having surface contamination that only affects Ta grown at high temperature seems unlikely.
@@
For the case of clean Ta films at dilution refrigerator temperatures (35mK in this work), thermally activated quasiparticles are exponentially suppressed Aumentado, Catelani, and Serniak (2023) and are not expected to be an observable source of MW loss in CPW resonators. Suppression of superconductivity at the Ta/sapphire interface could in principle result in an equilibrium quasipartile density responsible for the observed loss, but the mechanism of suppression is not clear. It has long been established that the gap energy of a BCS superconductor in contact with a trivial insulator is only negligibly reducedDeGennes (1964). Even normal metals such as gold (Au) tend to proximitize and high performing microwave devices have been demonstrated using Au capped Nb and TaBal et al. (2024); Chen et al. (2024). For these reasons, trivial suppression of superconductivty at the interface to the point that the observed MW loss may be explained by equilibrium quasiparticles seems unlikely.
@@ Acknowledgements.
All authors declare no financial or non-financial competing interests. The datasets used and/or analyzed during the current study available from the corresponding author on reasonable request. This work is a contribution of the U.S. government and is not subject to U.S. copyright.

### v2

@@ III.3 Modifying the Ta/ interface
For the second test, referred to as Sample C, the Al2O3 (0001) substrate was treated with an argon (Ar) plasma in the growth chamber in-situ at room temperature before the high vacuum outgas and growth of Ta directly on the substrate at 630. The plasma exposure was performed at 10mTorr using an RF power of 50W for 120 seconds. This plasma condition was found to etch aluminum at 1.6 nm/min in a separate calibration that is provided here for repeatability. The outgas temperature (880) used is not expected to result in restructuring of the Al2O3(0001) surface, which occurs at temperatures above 1000 Wildes, Mayer, and Theis-Brol (2001).This plasma treatment induces damage to the sapphire surface, and the resulting Ta film was found to be polycrystalline and textured in XRD measurements. The damage to the sapphire surface was observed with RHEED and more details are included in Appendix B.
@@
Figure 4(b) shows higher resolution scans of the -Ta(222) reflections from which the difference in the out-of-plane lattice constant of the Ta films grown with and without the Nb inter-layer is computed to be less than 1%. Fig. 4(c) shows - scans (in-plane sample rotation) of non-specular substrate and film peaks of Sample A, which confirm that the Ta is epitaxially arranged to the substrate having the same rotational relationship stated earlier for the epitaxial Nb. Non-specular XRD measurements of Sample C were also performed which showed in-plane texturing. Additional XRD data is included in Appendix C.
Figure 5 includes atomic force micrographs taken from Samples A, B, and C. For Samples A and B, (the epitaxial Ta with and without the 5nm thick Nb interlayer), the surface morphology was smooth and shapeless with little difference between the two. For Sample C (the Ta film grown on plasma-treated substrate), elongated, grain-like features were observed, consistent with previous reports of textured Ta and Nb films Clavero et al. (2012); Jones et al. (2023). The direction of the elongation reflects the in-plane texturing of the film where the same 6 rotational domains measured with XRD are observed. RMS roughness values extracted from AFM data is included in Table 2 for these samples.
Figure 6 shows temperature dependent resistivity measurements of Samples A, B, and C. For the epi--Ta(111) (Samples A and B), the 5nm Nb interlayer had little effect on film resistance, RRR, or . The RRR of the textured Ta(110) film grown on the plasma damaged sapphire (Sample C) was lower, reflecting less structural order in the polycrystalline film. The of all -Ta films was between 4.2-4.5K and the complete data is included in Table 2.
@@ IV Discussion
Consider Samples A, B, and C as indicated in Table 2. All three films were grown on c-plane sapphire at the same temperature of 630: Sample A is epi--Ta(111) grown directly on sapphire, Sample B is epi--Ta(111) grown on a 5nm thick Nb interlayer on sapphire, and Sample C is textured--Ta(110) grown on plasma treated sapphire. The of resonators made from Sample A are markedly lower than B and C. The microscopic mechanism of this loss is of interest, and some potential sources are discussed here in light of consistency with the results presented in this work.
@@
Initially, the origin of the observed interface loss was thought to be either due to reactions between the Ta and the Al2O3 substrate or chemical contamination on the sapphire surface before growth, but neither explanation is fully consistent with the results. First addressing interface contamination: the substrates used for all samples presented in this study (with exception of Sample C) were prepared in the same way and thus had very similar levels of surface contamination before film growth. High was observed for all Nb films grown, and also for the Ta films grown at lower temperature. Having surface contamination that only affects Ta grown at high temperature seems unlikely.
@@
For the case of clean Ta films at dilution refrigerator temperatures (35mK in this work), thermally activated quasiparticles are exponentially suppressed Aumentado, Catelani, and Serniak (2023) and are not expected to be an observable source of MW loss in CPW resonators. Suppression of superconductivity at the Ta/sapphire interface could in principle result in an equilibrium quasiparticle density responsible for the observed loss, but the mechanism of suppression is not clear. It has long been established that the gap energy of a BCS superconductor in contact with a trivial insulator is only negligibly reducedDeGennes (1964). Even normal metals such as gold (Au) tend to proximitize and high performing microwave devices have been demonstrated using Au capped Nb and TaBal et al. (2024); Chen et al. (2024a). For these reasons, trivial suppression of superconductivty at the interface to the point that the observed MW loss may be explained by equilibrium quasiparticles seems unlikely.
@@
The observed differences in microwave loss with interface structure may be caused by differences in oxygen bonding across the Ta/Sapphire interface. For future work, In-situ RHEED and X-ray photoelectron spectroscopy (XPS) studies may be used to study the chemical bonding between tantalum and oxygen in the first few monolayers of growth as the interface is formedFujii et al. (1996); Beran and Masek (2019). Lattice-site specific bonding of tantalum to oxygen may be studied using X-ray standing wave (XSW) techniquesChen et al. (2024b) which may prove insightful comparing the ordered, epitaxial interface of Sample A to the disordered interface of Sample C, for example. Notably, all films discussed in this work were grown on a sapphire surface that underwent a high temperature anneal in vacuum in order to obtain the same surface chemistry and Al-O stoichiometry. The oxygen content of this surface may be experimentally controlled by ion bombardment, oxygen dosing, or thermal annealing for exampleGillet and Ealet (1992) which may correlate to microwave loss.
@@ Acknowledgements.
All authors declare no financial or non-financial competing interests. The datasets used and/or analyzed during the current study available from the corresponding author on reasonable request. This work is a contribution of the U.S. government and is not subject to U.S. copyright. Certain commercial equipment, instruments, or materials are identified in this paper to foster understanding. Such identification does not imply recommendation or endorsement by the National Institute of Standards and Technology, nor does it imply that the materials or equipment identified are necessarily the best available for the purpose.
Appendix A: Additional Epitaxial Ta Films
The most important finding in this work, namely that the highest quality epitaxial Ta films yield lossy resonators when grown directly on sapphire but are high performing when grown on a thin, epitaxial Nb inter-layer has been reproduced in samples not included in the main text. Table 3 includes transport data for 4 additional epitaxial -Ta samples along with those already included in the main text. The two additional Ta films grown directly on sapphire yielded resonators with , while the films grown on an epitaxial Nb interlayer yield resonators with , consistent with the conclusions outlined in the main text.
The crystalline structure of these epitaxial -Ta films is very similar, and all test resonators underwent the same BOE cleaning procedure before measurement, however there are some important differences that should be emphasized. Sapphire wafers having both double- and single-side polishing (DSP/SSP) were used as well as several different vendors. In addition to any inherent differences between the substrates, the backside polish affects the wafer mounting configuration in the growth chamber and thus the surface growth temperature as well.
Samples having two additional surface preparations before film growth are also included; one with a chemical clean in stabilized Piranha solution held at for 10 minutes, and another annealed in air at in order to obtain a terraced sapphire surfaceDwikusuma, Saulys, and Kuech (2002); Vlasov et al. (2016) which was confirmed using AFM. All samples were annealed and outgassed at before cooling to the indicated growth temperature for film deposition.
Appendix B: RHEED characterization of the damaged sapphire surface
Ar plasma exposure is expected to both etch the sapphire substrate and induce structural damage to the sapphire surface. The structural damage to the starting surface of Sample C was observed using in-situ RHEED, where the bright diffraction pattern of the sapphire observed initially became dim (though still observable) after plasma exposure, and was accompanied by an increase in the diffuse background, indicating a less ordered surface. This diffuse RHEED pattern did not improve following high vacuum annealing at for 2 hours and dropping to the growth temperature of . RHEED patterns of this sample after Ta growth were consistent with a textured film having both out-of-plane and in-plane texture. This texture was confirmed with XRD after the sample was cooled and removed from the growth chamber. Figure 8 shows in-situ RHEED images of the starting sapphire surfaces of Samples A and C, the damaged sapphire surface from Sample C, and the resulting 100nm thick Ta films.
Appendix C: XRD pole figures
In addition to the non-specular measurements included in Fig. 4(c), pole figure measurements were also performed using the same diffractometer configuration outlined in the Methods section of the main text. These data complement the specular () measurements included in Figs. 1 and 4 and are used here to examine epitaxy and in-plane texturing. Figure 9 shows polar plots of the Ta/Nb(110) reflections (2) for three samples included in the main text: the epitaxial Ta film of Sample A, the polycrystalline textured film of Sample C, and the Nb film grown at room temperature.
Turning first to Sample A, three distinct peaks are observed separated by a polar angle of corresponding to the Ta(110), (101), and (011) crystal planes of the (111) oriented Ta film. This measurement confirms that the film is epitaxial with a single crystalline orientation. Turning next to Sample C, a clear peak is observed at zero azimuthal angle, consistent with specular measurements which had shown the film is (110) oriented. The off-specular peaks from this film are observed to have some structure, with peak maxima slightly misaligned from the sapphire high symmetry directions which confirms this film is highly textured out of plane in the (110) orientation with some in-plane texturing as well. The Nb film grown at room temperature has similarity to both Samples A and C, containing both epitaxial (111) grains and (110) grains having similar in-plane texture to Sample C.
Appendix D: AFM measurements of growth temperature series
Atomic force microscopy was performed on all films presented in this study. Figure 10 shows AFM images of the Nb and Ta films grown at varying substrate temperature. At higher growth temperatures, triangular features are observed in the surface morphology. For Nb, the triangular morphology was observed for both the samples grown at and . For Ta, triangular morphology is also observed, though the required temperature is higher than that for Nb.
Appendix E: Quality factor measurement details and power dependence
The procedure used for measuring internal quality factor and estimating photon occupancy of the hanger resonators is as follows: first, the individual resonators are mapped out manually using a vector network analyzer (VNA) to obtain the approximate resonance frequency and needed frequency span to encompass the resonance line-shape. Transmission (S21) spectra are then acquired over a range of VNA output powers. These data are then fit to the diameter correction method modelKhalil et al. (2012), which for hanger resonators reduces toMcRae et al. (2020):
In this expression, , is the total quality factor defined by , is the resonance frequency, and are the internal and external quality factors respectively, and is the correction angle representing asymmetry in the Lorenzian lineshape, More details on the derivation of these expressions and physical origin of the correction angle is provided by Khalil et al. (2012); Rieger et al. (2023).
The power used to probe the resonators may be expressed as an average photon number in the cavity. For a hanger resonator driven on resonance, this photon number is given by the following expression Bruno et al. (2015).
where is Planck’s constant and is the incident power at the device. The parameters , , and are extracted from the fit to the DCM model above. is estimated using the power output of the VNA having a dynamic range of +20 dBm to -87 dBm, with the known attention of 80 dB from the VNA to the mixing chamber of the dilution refrigerator. This includes 70 dB of attenuation from cryogenic attenuators and an estimated 10 dB of cabling losses.
The power dependence of resonator is commonly used to extract the contribution of saturable loss sources or two-level systems (TLS) which have been suggested to be a source of decoherence in state-of-the-art superconducting qubitsMcRae et al. (2020). Figure 11 shows resonator vs estimated photon number for several chips that were presented in this work as well as some example traces and fits to the DCM model. The data included in Fig.11(a) was simply selected in order to provide representative curves while avoiding overlapping data sets. Throughout this work, a total of 16 chips were measured in this way, though data for only 4 chips is included in this plot for clarity. For the best performing resonators (epitaxial Ta grown on a Nb nucleation layer), varies by nearly two orders of magnitude in the range studied demonstrating the contribution of TLS to the low power loss. In contrast, of the worst performing resonators (epitaxial Ta grown directly on Al2O3(0001)) show a weak or negligible power dependence in the range accessible.
| Substrate/Polish. | Surface Prep | Film | Growth Temp. | RRR | () | Resonator mean |
| () | (Low-power) | |||||
| Vendor A, SSP (Sample A)111Included in main text | Solvent Clean/HV Anneal | 100nm Ta/ | 630 | 47 | 4.35 | 8.3k |
| Vendor A, SSP11footnotemark: 1 | Solvent Clean/HV Anneal | 100nm Ta/ | 520 | 29 | 4.24 | 7.4k |
| Vendor B, SSP | Stabilized Piranha/HV Anneal | 100nm Ta/ | 800 | - | - | < 20k |
| Vendor C, DSP222Wafer underwent additional Josephson junction processing steps | 3 hr.@1100C in Air/HV Anneal | 200nm Ta/ | 750 | - | - | 32k |
| Vendor A, SSP (Sample B)11footnotemark: 1 | Solvent Clean/HV Anneal | 95nm Ta/5nm Nb/ | 630 | 45 | 4.35 | 520k |
| Vendor A, DSP | Solvent Clean/HV Anneal | 195nm Ta/5nm Nb/ | 645 | 78 | 4.33 | 1.3M |
| Vendor A, DSP22footnotemark: 2 | Solvent Clean/HV Anneal | 195nm Ta/5nm Nb/ | 645 | 52 | 4.29 | 740k |